Moralofthestory:

How many ounces of trail mix are in a bag that weighs 0.681 kilograms? Input only numeric values. (1 pound = 0.454 kg and 1 pound = 16 ounces)

itsme:

@supie ik i tag u a lot but dyk the answer?

Mercury:

\[\frac{ 0.681~kg }{ }\frac{ 1~lb }{ 0.454~kg }\frac{ 16~oz }{ 1~lb }\] noticehow the units cancel out and you're left with oz after doing the calculations